Method Substitution
5. Find the point of intersection of each pair of lines.
b) 4x+2y=7
-x-y=6
Please, i dont know how to do it thank you very much!!!!

Answer by stanbon(75887)

Find the point of intersection of each pair of lines.
b) 4x+2y=7
-x-y=6
----------------
Leave the 1st equation along;
Multiply the 2nd equation by 4.
----------------------------------
4x + 2y = 7
-4x- 4y = 24
------------------
Add the two equations to eliminate the x-terms:
-2y = 31
y = -31/2
---------------
Substitute into -x-y = 6 to solve for "x":
-x -(-31/2) = 6
x + 31/2 = 12/2
x = -19/2
==========================
Cheers,
Stan H.